IBM stock currently sells for 49 dollars per share. Over 12 months the price will either go by 11.5 percent or down by -7.0 percent. The risk-free rate of interest is 4.5 percent continuously compounded. What is the value of a call option with strike price 51 and maturity 12 months?

a.

0.12291

b.

1.9353

c.

2.1795

d.

1.3285
Answer #1

Solution:

Sl.No.

Particulars

Notation

Value

1

Spot Price

SP0

$ 49.00

2

Strike Price

EP

$ 51.00

3

Expected future Spot price – Lower Limit - FP1

( $ 49 * 93 % ) = $ 45.570

FP1

$ 45.570

4

Expected future Spot price – Upper Limit FP2

( $ 49 * 111.5 % ) = $ 54.635

FP2

$ 54.635

5

Value of call at lower limit

[ Action = Lapse, Since FP1 < EP. Therefore value = Nil ]

Cd

NIL

6

Value of call at upper limit

[ Action = Exercise, Since FP2 > EP. Therefore value = ( $ 54.635 - $ 51.00 = $ 3.635 ) ]

Cu

$ 3.635

7

Weight for the lower scenario

[FP1 / SP0 ] = ( 45.570 / 49 ) =

d

0.930

8

Weight for the upper scenario

[FP2 / SP0 ] = ( 54.635 / 49 ) =

u

1.115

9

Risk free rate of Return

r

0.045

10

Duration of the call

t

1 Year

11

Future value factor (Continuous Compounding factor) = er * t = e0.045 * 1 = e0.045 = 2.71828 ( 0.045) = 1.046028

f

1.046028

Note : The value of 2.71828 ( 0.045) is calculated using the excel formula =POWER(Number,Power)

=POWER(2.71828,0.045) = 1.046028

As per the Binomial Option Pricing formula the value of a call is given by the following formula:

Value of a Call = [ ( Cu * [(f-d)/(u-d) ] ) + ( Cd * [ (u-f)/(u-d) ] ) ] / f

Therefore applying the values from the table above to the formula we now have:

= [ ( 3.635*[ (1.046028 - 0.930 )/(1.115 – 0.930) ] ) + ( 0 *[ (1.115 – 1.046028 )/( 1.115 – 0.930) ] ) ] / 1.046028

= [ ( 3.635* [ (0.116028)/( 0.185 ) ] ] / 1.046028

= [ 3.635 * 0.627178 ] / 1.046028

= 2.279793 / 1.046028

= 2.179476

= 2.1795 ( when rounded off to four decimal places )

The value of a call option with strike price 51 and maturity 12 months = $ 2.1795

Thus the solution is Option c = $ 2.1795